Evaluate the following expression when x = 5 and z = 3.

z/4x

21 Fav MATH 0530-602 (6) Homework: Competency Builder 2 Score: 0 of 1 pt 8.2.51 Evaluate the following expression when x 5 and z # 2. The answer isType an integer or a fraction. Simplify your answer.) Enter your answer in the answer box and then click Check Answer All parts showing
Answer #1

Here we are required to find the value of the expression given by z/(4x) when z=2 and x=5,

so the evaluated value of the expression is= 2/(4*5)=1/10=0.1